# High-Coherence Kerr-cat qubit in 2D architecture

### v1

@@ VI Conclusion
Our results demonstrate a high-coherence Kerr-cat qubit in a scalable 2D architecture with high-fidelity single-qubit operations and minimal heating. This demonstration is enabled by integrating a wide band-block filter for efficient energy delivery, enabling strong light-matter coupling with negligible Purcell effect. We report improvement in the bit-flip time up to cats of size $\alpha^{2}=10$ with lifetimes higher than 1 ms in good agreement with our theoretical modeling of the circuit, including multi-photon processes. Our qubit was based on a SNAILmon oscillator with a single photon decay time $T_{1}=38.5$ $\mu s$.
Operating at a Kerr nonlinearity of 1.2 MHz, we get an energy gap $4K\alpha^{2}=$ 9.6-48 MHz for cats of size 2-10, which enables fast operation compared to the decay rates. For those device parameters, we demonstrate quantum non-demolition readout with QNDness of 98.3% for a cat of size 4 and 99.6% for a cat of size 8. We also introduce a resource-efficient way to implement the $X(\pi/2)$ gate with phase modulation of the stabilization drive. Combined with continuous and fast $Z(\theta)$ rotations, this is sufficient to have universal control over the Kerr-cat qubit. For a cat of size 4, we implement single qubit $Z(\pi/2)$ in 120 ns and the $X(\pi/2)$ in 320 ns with gate fidelities of $91.7\%$ and $90.8\%$ respectively, limited by state preparation since for the identity, we measure process fidelity of $92.7\%$ using process tomography. A detailed analysis of those gates in terms of noise bias and the removal of spam errors is the subject of future work.
With the goal of implementing quantum error correction codes with many physical Kerr-cat qubits the problem of finding an efficient pumping scheme that does not heat the cryogenic environment is of most importance. Our solution relied on integrating an effective band-block filter to remove the trade-off between strong capacitive coupling to the qubit and large Purcell decay to enable strong light-matter interactions. With 30 dB of isolation at the qubit frequency, large bandwidth, and negligible attenuation at the pump frequencies, this design has the potential to realize multi-qubit processors.
In conclusion, we introduced a novel scheme for efficient power delivery in 2D architecture to realize the Kerr-cat qubit with high coherence, fast logic gates, and state of art QND readout. This work extends the reach of the Kerr-cat qubit to strongly coupled multi-qubit 2D processors, which can be used to examine the fidelity of the noise-bias-preserving two-qubit CZ and CX gates [27] and quantum error correction schemes tailored for noise-biased qubits with the Kerr cats [38, 36, 37]. Future single-qubit experiments can focus on increasing the Kerr nonlinearity of the qubit without jeopardizing the lifetime to enable faster single-qubit gates with improved fidelity and examine the root causes of the saturation in the lifetime with controlled dissipation.
@@ VII Acknowledgments
We thank Chuan-Hong Liu, Akel Hashim, and Ravi Naik for valuable discussions. This work was supported by the U. S. Army Research Office under grant W911NF-22-1-0258
@@ A.2 Cat Quadrature Readout
Figure 5: Lifetime statistics. The average lifetime for multiple states is examined over the span of two hours with 100 iterations. The lifetime of the first excited state in the Fock basis |$\ket{n=1}$ can be used to estimate the single photon loss at the qubit frequency. The long-lived coherent states |$\ket{\alpha^{2}=\epsilon_{2}/K}$ also fluctuate over time, with an average lifetime of $601$ $\mu s$ for a cat of size 4 and $872$ $\mu s$ for a cat of size 8.
@@ Appendix B Master equation calculations
In order to simulate the lifetime of the coherent states, we consider the system to be linearly coupled to a macroscopic environment composed of a bath of linear resonators with continuous modes. The system environment coupling Hamiltonian in the rotating frame takes the following form
@@
where $V_{k}$ gives the coupling strength and $\hat{b}_{k}(\hat{b}_{k}^{\dagger})$ are the annihilation (creation) operators of an excitation of energy $\hbar\omega_{k}$ in the environment. Following Ref. 64, we obtain following master equation up to ${\cal O}(\varphi_{\rm zps}^{2})$,
@@ Appendix C Universal single-qubit gates for the Kerr cat qubit
A universal gate set for the Kerr cat qubit has been reported in the literature [20] based on continuous $Z(\theta)$ rotations and a discrete $X(\pi/2)$ rotation on the Bloch sphere. However, the discrete $X(\pi/2)$ relies on turning off the stabilization drive, which protects this qubit and the performance of this gate degrades very quickly with increasing cat size $\left|\alpha\right|^{2}$[32]. Therefore, it is necessary to realize an alternative approach with higher fidelity. Here, we describe the implementation of gates we introduced in the main text.
@@ C.1 Continuous Z($\theta$) rotation based on the quantum Zeno effect
Figure 6: Simulation of the detuning gate. Two $X(\pi/2)$ pulses are used to flip the value of $\innerproduct{Z}{Z}$, similar to the experimental sequence in Fig. 3(b). Sweeping the gate time $T_{g}$ and the maximum detuning $\delta_{0}$ according to Eq.42 shows multiple regions where the potential deformation leads to coherent tunneling between the energy wells.
@@ Appendix H Microwave control hardware
To ensure the thermalization of the qubit, we attenuate the lines heavily and use a gold-plated cryogenic attenuators from Quantum Microwave at the mixing chamber in combination with eccosorb filters. To avoid dumping heat at the mixing chamber (MXC), we use directional couplers to attenuate the signal at the MXC plate while directing the signal to the 4 K with 50 ohm terminations. The output signal is first amplified with a traveling wave parametric amplifier (TWPA) near the quantum limit with a gain of around 20 dB at the mixing chamber [70], then a HEMT amplifier at 4 K, and a low-noise amplifier at room temperature.

### v2

@@ VI Conclusion
Our results demonstrate a high-coherence Kerr-cat qubit in a scalable 2D architecture with high-fidelity single-qubit operations and minimal heating. This demonstration is enabled by integrating a wide band-block filter for efficient energy delivery, enabling strong light-matter coupling with negligible Purcell effect. We report improvement in the bit-flip time up to cats of size $\alpha^{2}=10$ with lifetimes higher than 1 ms, in good agreement with our theoretical modeling of the circuit. Our qubit was based on a SNAILmon oscillator with a single photon decay time $T_{1}=38.5$ $\mu s$.
Operating at a Kerr nonlinearity of 1.2 MHz, we get an energy gap $4K\alpha^{2}=$ 9.6-48 MHz for cats of size 2-10, which enables fast operation compared to the decay rates. For those device parameters, we demonstrate quantum non-demolition readout with QNDness of 98.3% for a cat of size 4 and 99.6% for a cat of size 8. We also introduce a resource-efficient way to implement the $X(\pi/2)$ gate with phase modulation of the stabilization drive. Combined with continuous and fast $Z(\theta)$ rotations, this is sufficient to have universal control over the Kerr-cat qubit. For a cat of size 4, we implement single qubit $Z(\pi/2)$ in 120 ns and the $X(\pi/2)$ in 320 ns with gate fidelities of $91.7\%$ and $90.8\%$ respectively, limited by state preparation using process tomography. A detailed analysis of those gates in terms of the noise bias without SPAM errors is the subject of future work.
With the goal of implementing quantum error correction codes with many physical Kerr-cat qubits the problem of finding an efficient pumping scheme that does not heat the cryogenic environment is of most importance. Our solution relied on integrating an effective band-block filter to remove the trade-off between strong capacitive coupling to the qubit and large Purcell decay to enable strong light-matter interactions. With 30 dB of isolation at the qubit frequency, large bandwidth, and negligible attenuation at the pump frequencies, this design enables the realization of multi-qubit processors.
In conclusion, we introduced a novel scheme for efficient power delivery in 2D architecture to realize the Kerr-cat qubit with high coherence, fast logic gates, and state of art QND readout. This work extends the reach of the Kerr-cat qubit to strongly coupled multi-qubit 2D processors, which can be used to examine the fidelity of the noise-bias-preserving two-qubit CZ and CX gates [30] and quantum error correction schemes tailored for noise-biased qubits with Kerr cats [41, 39, 40]. Future single-qubit experiments can focus on increasing the Kerr nonlinearity of the qubit without jeopardizing the lifetime to enable faster single-qubit gates with improved fidelity and examine the root causes of the saturation in the lifetime with controlled dissipation.
@@ VII Acknowledgments
We thank Chuan-Hong Liu, Akel Hashim, and Ravi Naik for valuable discussions on the experiment. We are grateful to Alexei Marchenkov and Alexander Tselikov for the blueprint of the on-chip RF filter. This work was supported by the U. S. Army Research Office under grant W911NF-22-1-0258.
@@ A.2 Cat Quadrature Readout
Figure 5: Lifetime statistics. The lifetime for multiple states is examined over the span of two hours with 100 iterations. The lifetime of the first excited state in the Fock basis |$\ket{n=1}$ can be used to estimate the single photon loss at the qubit frequency. The long-lived coherent states |$\ket{\alpha^{2}=\epsilon_{2}/K}$ also fluctuate over time, with an average lifetime of $601$ $\mu s$ for a cat of size 4 and $872$ $\mu s$ for a cat of size 8.
@@ Appendix B Master equation calculations
In order to simulate the lifetime of the coherent states, we consider the system to be linearly coupled to a macroscopic environment composed of a bath of linear oscillators with continuous modes. The system environment coupling Hamiltonian in the rotating is:
@@
where $V_{k}$ gives the coupling strength and $\hat{b}_{k}(\hat{b}_{k}^{\dagger})$ are the annihilation (creation) operators of an excitation of energy $\hbar\omega_{k}$ in the environment. Following Ref. 68, we obtain following master equation up to ${\cal O}(\varphi_{\rm zps}^{2})$,
@@ Appendix C Universal single-qubit gates for the Kerr cat qubit
A universal gate set for the Kerr cat qubit has been reported in the literature [20] based on continuous $Z(\theta)$ rotations and a discrete $X(\pi/2)$ rotation on the Bloch sphere. However, the discrete $X(\pi/2)$ relies on turning off the stabilization drive, which protects this qubit and the performance of this gate degrades very quickly with increasing cat size $\left|\alpha\right|^{2}$[35]. Therefore, it is necessary to realize an alternative approach with higher fidelity. Here, we describe the implementation of the gates we introduced in the main text.
@@ C.1 Continuous Z($\theta$) rotation based on the quantum Zeno effect
Figure 6: Simulation of the detuning gate. Two $X(\pi/2)$ pulses are used to flip the value of $Z\right>$, similar to the experimental sequence in Fig. 3(b). Sweeping the gate time $T_{g}$ and the maximum detuning $\delta_{0}$ according to Eq.42 shows multiple regions where the potential deformation leads to coherent tunneling between the energy wells.
@@ Appendix H Microwave control hardware
To ensure the thermalization of the qubit, we attenuate the lines heavily and use a gold-plated cryogenic attenuator from Quantum Microwave at the mixing chamber in combination with eccosorb filters. To avoid dumping heat at the mixing chamber (MXC), we use directional couplers to attenuate the signal at the MXC plate while directing the signal to the 4 K with 50-ohm termination. The output signal is first amplified with a traveling wave parametric amplifier (TWPA) with a gain of around 20 dB at the mixing chamber [74], then a HEMT amplifier at 4 K, and a low-noise amplifier at room temperature. The signal is then downconverted with an IQ mixer, filtered around the demodulation frequency, and amplified before sending it to the ADC.